OpenStudy (anonymous):

b/-3+5=7 A. b=-2/3 B. b=-4 C. b=-36 D. b=-6

OpenStudy (jhannybean):

\[\frac{ b }{ -3+5 }=7\] so what does -3 + 5 =?

OpenStudy (jhannybean):

First solve that. Then we will move onto the next step :)

OpenStudy (zzr0ck3r):

I think its (b/-3) + 5 = 7

OpenStudy (jhannybean):

oh -________________________- parenthesis would help sooooo much.

OpenStudy (zzr0ck3r):

yeah I don't know for sure...

OpenStudy (jhannybean):

Me neither....it could be either way.....

OpenStudy (zzr0ck3r):

just going by order of operations/what a computer would read from it.

OpenStudy (zzr0ck3r):

or calculator...

OpenStudy (jhannybean):

if it was \(\large -(\frac{ b }{ 3 })+5 = 7\) the person would still need to know what (7-5) equals... and he/she is not answering my question....

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

Ah sorry I didn't know you guys answered so fast. -3+5 is 2 so...

OpenStudy (zzr0ck3r):

what is the original question? (b/-3) + 5 = 7 or b/(-3+5) = 7

OpenStudy (jhannybean):

^ ?

OpenStudy (anonymous):

(b/-3)+5=7

OpenStudy (jhannybean):

Darn it @zzr0ck3r Nowi owe him $5.thanks @Zoobidoo :P jk.

OpenStudy (zzr0ck3r):

subtract 5 from both sides, then multiply by -3 on both sides.

OpenStudy (zzr0ck3r):

you are picking the thing apart so that you are left with b on one side and everything else on the other, so first get rid of the 5(subtract 5) then get rid of the -3(multiply by -3)

OpenStudy (jhannybean):

\[-(\frac{ b }{ 3 })+5 = 7\] so now you have to evaluate b, that means taking all the numbers and putting them on the other side of b, so you get \[-(\frac{ b }{ 3 })= 2\] and then multiply both sides by (-3) to get b

OpenStudy (anonymous):

ok got it!

OpenStudy (zzr0ck3r):

whats the answer?

OpenStudy (anonymous):

im working on it...

OpenStudy (jhannybean):

What do you get when you mutiply (-3 * 2)?

OpenStudy (anonymous):

-6

OpenStudy (jhannybean):

goood job!!

OpenStudy (anonymous):

Thank you guys so much!

OpenStudy (jhannybean):

no problem!!